how do you find the missing assets for Company C and Required information [The following information applies to the questions displayed below.) The following financial statementthe Stock Issuances for Company D and Liabilities for Company E
Answer #1

a) Company C

Beginning stockholder's equity = 26240-14169 = 12071

Ending Stockholder's equity = 12071+9750+5200-5875 = 21146

Total Assets on December 31,2019 = 21146+15101 = 36247

b) Company D :

Beginning equity = 72980-50356 = 22624

Ending equity = 83720-40185 = 43535

Stock issuance = 22624+13911-43535 = 7000

c) Company E :

Ending equity = 126960-100298 = 26662

Beginning equity = 26662-6500-8776+11000 = 22386

Beginning liabilities = 111930-22386 = 89544